# Uncovering Autoregressive LLM Knowledge of Thematic Fit in Event Representation

### v1

@@ 3.3 Models and Hardware
We conducted the experiments using two models, the latest as of the time of running the experiments. (1) OpenAI GPT-4-turbo (gpt-4-0125-preview, hereafter ‘GPT’) (Achiam et al., 2023) as a closed model, and (2) Meta CodeLlama2 (codeLlama2-70B-instruct, hereafter ‘Llama’) (Roziere et al., 2023) as an open model (even if not fully open). See details and required hardware in Appendix §F.
3.4 Experimental Settings
We conducted a preliminary study to determine good temperature and top_p parameter values for the LLMs. We used GPT-3.5 on a subset of the Fer-Ins dataset (20 samples $\sim$ 8%) and ran a parameter sweep over temperature values (0.0, 0.5, 0.9) and top_p values (0.5, 0.7, 0.95). Temperature values were chosen near-uniformly across the scale to examine a wide range of the model creativity setting. Higher values correspond to higher randomness (aka "creativity") in the output. High values of top_p limit the randomness in the output to more probable tokens. Since we were more interested in accurate output, and less interested in creative writing for our purposes here, we tested top-p threshold values on the higher half of the scale. We then sort the human rating list and the generated scores list and compute a Spearman’s Rank Correlation Coefficient ($\rho$) between these two sorted lists. Temperature 0.0 and top_p 0.95 showed the highest correlation coefficient with human judgments on our sample. We set the max_tokens parameter based on the length of the prompt and its expected output. For experiments with Head Lemma Tuples input, we set max_tokens to 100, whereas for Generated Sentences input, we set it to 300. As Step-by-Step Prompting  is expected to be lengthy to clarify the underlying reasons, we set max_tokens to 600 there. We used these settings in all experiments. [^3] The source code is available at: https://github.com/SafeyahShemali/LLM-Thematic-Fit-24
@@ 4 Results
This section consists of the following parts. §4.1: the GPT  results. §4.2: the Llama results.
4.1 GPT Results
Experiment 1.1, the simplest in settings and the most similar to the baselines (using head lemmas), yielded scores that are pronouncedly higher than the baselines across the board. [^4] Except for baseline B3c on Pado, whose results are slightly higher than GPT’s. But B3c reports the highest scores in several runs, while the average is reported in B3b, which is slightly lower than the score in Exp.1.1. See Table 2. This is a surprising and impressive achievement for GPT, since to the best of our knowledge, it does not have an explicit self-reflection capability nor was it trained on thematic fit labels.
@@ 4.1 GPT Results
Step-by-Step Prompting  dramatically improved Ferretti’s Location (Fer-Loc) results compared to Simple Prompting, yielding a new state-of-the-art (.31 absolute improvement over the strongest baseline). Results in all Step-by-Step Prompting  experiments (Exp.3.x or 4.x) were higher than all baselines on the Ferretti datasets.
For Fer-Ins, where the highest score with Step-by-Step Prompting was .68, the correlation coefficients improved with Step-by-Step Prompting for .18-.20 (absolute) over the strongest baseline. However, the maximum score on Fer-Ins (.74) was obtained with Simple Prompting  and categorical output form (Exp.1.2).
Surprisingly, Step-by-Step Prompting  achieved much lower scores on Pado and McRae compared with Simple Prompting in Exp.3.x and slightly lowered scores in Exp.4.x. Overall, there was a sizable gain in Fer-Loc, but only there.
Results from Exp.1, 2, and 4 were generally higher than the baseline results. [^5] But for Pado, see footnote 4. In Exp.3.x, the scores on Pado were lower than all of the baselines. Similarly, the scores on McRae were lower than most baselines. However, the corresponding scores in Exp.4.x (Step-by-Step Prompting  with generated sentences) were much higher.
@@
Using Generated Sentences instead of Lemma Tuples yielded large gains with Step-by-Step Prompting  on Pado and McRae. The gain in Pado was .26 (absolute) with a numeric score and .25 (absolute) with a categorical score. The gain in McRae was .09 (absolute) with either a numeric or categorical score.
Using Generated Sentences instead of Lemma Tuple with Simple Prompting  lowered thematic fit scores on all test sets except McRae in Exp.2.2 vs. Exp.1.2 (Simple Prompting  and categorical scores), where we got a slight gain, if any, of .03 (absolute).
@@
Using predefined categorical output compared to numeric output (Exp.x.2 vs. Exp.x.1) raised the correlation coefficient across the board in all experiments and test sets, with few exceptions. Most notable of which were the lower scores of Exp.1.2 compared with 1.1 on Pado and McRae. The other case was with Pado in Exp.4.2.

### v2

@@ 3.4 Models and Hardware
We conducted the experiments using closed and open models, the latest available at the time of running the experiments. Closed models (as they are often best performing, but less reproducible): OpenAI gpt-4-0125-preview (achiam2023gpt; openai2023gpt4turbo) with unconfirmed 1.76 trillion parameters (hereafter ‘GPT4-Turbo’) and gpt-4.1-2025-04-14 (openai2025gpt41docs; OpenAI_GPT-4.1_2025) with 1.8 trillion parameter (hereafter ‘GPT4.1’). Open models (more controlled and reproducible, even if not fully open): llama3.2:3b-instruct-q4_K_M (Dubey2024Llama3; meta2024llama), a quantized version of Meta Llama3.2 with 3 billion parameter (hereafter ‘Llama’) and qwen2.5:7b-instruct-q4_K_M (Yang2024Qwen2.5) with 7.61 billion parameters (hereafter ‘Qwen’). Closed model inference utilized the OpenAI API. Open model inference utilized Ollama (Ollama2023), as it is resource-efficient.
@@ 4 Results
Overall, our results set new state of the art for all four tasks, especially with GPT4-Turbo using Step-by-Step Prompting (except for Pado, where GPT4-Turbo tied with the highest baseline). See Table 2 and details below.
4.1 Closed Models
Both closed models (GPT4-Turbo and GPT4.1), when using Step-by-Step Prompting, outperformed the baselines by a large margin, except for Pado (where GPT4-Turbo tied with the highest baseline, and GPT4.1 was weaker than all baselines). With respect to the three axes, the models’ performance differed as follows:
@@ 4.1 Closed Models
GPT4-Turbo demonstrated superior performance across the board and among all models, including GPT4.1, see Table 3. Step-by-Step Prompting experiments (Exp.3.x and 4.x) yielded the best scores, where absolute improvement over the strongest baseline was .23 for Fer-Loc, .22 for Fer-Ins, .23 for McRae, and a tie with Pado. Simple Prompting also outdid the baselines (except Pado).
GPT4.1 Step-by-Step Prompting also outdid all baselines except Pado, but curiously lower than GPT4-Turbo, in spite of being newer. Simple Prompting gave mixed results w.r.t. the baselines.
@@
Using Generated Sentences did worse or similar to using Lemma Tuples in all tasks and LLMs, except Pado with GPT4-Turbo.
@@
Using predefined categories compared to numeric output (Exp.x.2 vs. Exp.x.1) had zero, or small mixed effect, contrary to our hypothesis.
With GPT4-Turbo, using predefined categorical output compared to numeric output (Exp.x.2 vs. Exp.x.1) leads to a mixed effect across the board. Regardless of input and reasoning forms, the categorical score is lower than the numerical score in Fer-Loc. For Fer-Ins and Pado, the categorical score is either the same or higher than the numerical score, as the gain in Fer-Ins was .04 in Exp2 and in Pado was .04 in Exp3. Using categorical output in McRae lowers the score in general, except for Exp4, where the gain was .02.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .63 .64 .35 .54
Exp.1.2 .61 .65 .40 .58
Exp.2.1 .48 .49 .42 .55
Exp.2.2 .46 .54 .45 .56
Exp.3.1 .69 .70 .48 .65
Exp.3.2 .68 .68 .46 .66
Exp.4.1 .57 .66 .58 .58
Exp.4.2 .58 .66 .59 .59
Table 2: GPT4-Turbo Spearman Rank Correlation ($\rho$). BG = greenberg2015improving, B0 = tilk2016event, B1 = hong2018learning, B2 = Marton2021ThematicFB, B3 = muthupari2022s as baselines for our experiments. Subscripts a-c denote specific models there (Glove-shared-not-tuned, etc.). Exp.m.nm.n = see Table 1. All $\rho$’s had p-values $10^{-13}$.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .51 .56 .40 .64
Exp.1.2 .50 .56 .40 .58
Exp.2.1 .32 .33 .33 .40
Exp.2.2 .26 .37 .33 .37
Exp.3.1 .59 .59 .45 .64
Exp.3.2 .56 .59 .42 .63
Exp.4.1 .56 .53 .42 .44
Exp.4.2 .53 .53 .42 .46
Table 3: GPT4.1 Spearman Rank Correlation ($\rho$). Legend is as in Table 2. All $\rho$’s had p-values $10^{-5}$.
4.2 Open Models
The performance of the open models was lower than the closed models, and lower than several baselines across the board, except for Fer-Loc using Step-by-Step Prompting with Llama and McRae with Qwen (mostly in Exp.4.x. Details are below:
Axis 1: Reasoning Form
Despite being mostly lower than the baselines, Step-by-Step Prompting has a positive effect in Fer-Loc and McRae. In Pado, the effect was detrimental with Llama (Table 4), but positive with Qwen (Table 5).
Axis 2: Input Form
Using Generated Sentences instead of Lemma Tuples had mostly no effect with Llama, except for gains in McRae. With Qwen there were mostly gains except for lower scores in Exp.4.x in FerLoc.
Axis 3: Output Form
Using predefined categories in the output, compared to numeric output (Exp.x.2 vs. x.1), yielded lower scores with Llama, except in Pado Exp.3.x and 4.x, but the correlation scores there are too low to trust these differences are meaningful. With Qwen  results were worse in Fer-Ins and McRae, and mixed elsewhere.
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .32 .33 .27 .16
Exp.1.2 .30 .23 .26 .13
Exp.2.1 .32 .33 .31 .25
Exp.2.2 .29 .22 .22 .23
Exp.3.1 .46 .29 .05* .23
Exp.3.2 .41 .24 .10* .20
Exp.4.1 .46 .28 .06* .27
Exp.4.2 .41 .24 .12* .21
Table 4: Llama Spearman Rank Correlation ($\rho$). Legend is as in Table 2. All $\rho$’s had p-values $10^{-3}$ except cells with * for which p-values were between (.01 - .22).
Model Dataset Fer-Loc Fer-Ins Pado McRae
BG: GSD2015 .29 .42 .53 .36
B0: NN RF .44 .45 .52 .38
B1: ResRoFA-MT .46 .48 .53 .43
B2: 20% subset v2 - - .43 .44
B3a: Glove-shared-not-tuned .30 .34 .49 .31
B3b: Glove-shared-tuned .29 .29 .53 .33
B3c: Glove-shared-tuned with 20%M .35 .43 .59 .43
Exp.1.1 .29 .43 .28 .42
Exp.1.2 .38 .37 .19 .34
Exp.2.1 .31 .44 .31 .48
Exp.2.2 .37 .38 .22 .39
Exp.3.1 .41 .36 .32 .40
Exp.3.2 .41 .32 .36 .35
Exp.4.1 .31 .37 .33 .51
Exp.4.2 .22 .36 .37 .47
Table 5: Qwen Spearman Rank Correlation ($\rho$). Legend is as in Table 2. All $\rho$’s had p-values $10^{-3}$.
5 Discussion
The core questions we aim to address in this work are: Do LLMs have the kind of linguistic knowledge required for thematic fit estimation? If so, (Q1) does Step-by-Step Prompting help utilize the model’s internal linguistics knowledge for this task? (Q2) Would providing more context by including sentences as part of the input help? (as opposed to providing only lemma tuples with semantic roles) (Q3) Do predefined thematic fit output categories better elicit the model’s internal linguistic knowledge compared to numeric output?
Overall, we conclude that closed models acquired much of the linguistic knowledge required for thematic fit estimation. Results with both GPT models on each of the test sets achieved or even surpassed all previous baselines, which were trained on labeled linguistic data (even if not directly trained for thematic fit estimation). This is impressive, [^9] The available inter-annotator agreement or correlation information for our testsets indicates the upper bound for these tasks is between the mid-60’s and mid-70’s. GPT4-Turbo’s scores got close, but no single setting was always the best. assuming closed models were not trained on this task (sadly, we cannot know for sure because they are, well, closed). Next, we discuss the findings by the three axes (as defined in §3) and then move to analyzing additional aspects.
(Q1) Axis 1: Reasoning Form
For GPT models (closed models), Step-by-Step Prompting fulfilled its intended goal of providing a “scratchpad” to help them calculate inferences given their internal knowledge (with the exception of GPT4.1 in Pado). In contrast, Step-by-Step Prompting didn’t help with open models (except for Fer-Loc with Llama and for McRae with Qwen).
Why was there a difference in the Step-by-Step Prompting effect between the closed and open models? Step-by-Step Prompting is designed to help the models reason by first eliciting useful properties of arguments and roles, before the final thematic fit estimation. However, our analysis on a sample of random tuples revealed that open models are weak at inferring the essential properties that are most expected for an argument, to serve well in a specific semantic role. Input from this step, containing invalid or irrelevant properties, often derailed the final reasoning output quality (See Appendix §E).
Why did GPT4-Turbo outperform GPT4.1 in spite of being older? We observed that GPT4.1’s ratings were often more ‘cautious’, meaning it rarely output very high or very low scores. We verified this by calculating the variance of each model’s output for each task and experimental setting (Appendix §B, Tables 7-8), finding a negative correlation between the variance and the correlation score ($\rho$). The lower variance of GPT4.1 means it is more likely to do worse on clear-cut inputs (with very high or very low thematic fit).
(Q2) Axis 2: Input Form
In principle, providing more context enhances models’ ability to comprehend complex tasks. This was generally reflected in our experiments with open models, but not closed models (with few exceptions).
Why do generated sentences only help open models? Although the open models did not outperform the closed models, they often better leveraged the input with generated sentences, and improved the correlation score ($\rho$) relative to the tuple input. But sentence input is likely to outperform tuple input only if these sentences are “good” (semantically coherent, with the given argument in the given role for the given predicate). Moreover, “bad” sentences are likely to derail the models’ performance. Therefore, we analyzed a random sample of 20 tuples, 5 generated sentences for each, altogether 100 sentences per model. We found that the open Qwen did better in filtering out “bad” sentences (in spite of having fewer parameters): its semantic filtering let through no False Positives (FPs), thus successfully leveraging the sentence inputs. In contrast, the closed GPT4-Turbo let through 6 FPs (Table 6). We saw similar results comparing filtering by Qwen vs. GPT4-Turbo, both on GPT4-Turbo’s sentences. [^10] Note that Qwen had higher False Negative rate, but this would only result in regressing to tuple input (§3.1 Axis 2). With it, closed models clearly generated more “good” sentences (80 vs. 60 True Positives + False Negatives in our sample). Therefore we caution against jumping into conclusions which model gained more relevant knowledge.
Acceptable Sentence? Qwen GPT4- Turbo Qwen filtering GPT4-Turbo
True Positive 18 52 26
True Negative 40 15 20
False Positive 0 6 1
False Negative 42 27 53
Table 6: Semantic filtering of “bad” generated sentences, random sample size: 100. “Bad” = True Negatives + False Positives. Left 2 columns: model both generated and filtered.
(Q3) Axis 3: Output Form
There are arguments in favor of either categorical and numeric output. On the one hand, we expect predefined output categories to yield better results than numeric output, because LLMs are primarily designed for text-based tasks. On the other hand, in spite of this, LLMs have also shown impressive performance in math (even if not perfect). Also, the ground-truth thematic fit scores were averaged across the participants, providing a graded scale, and therefore perhaps a numeric output would better correlate with that. Indeed, output form had mostly small-to-zero effect (up to 0.02 difference) for closed models, but categorical output gave mixed results, mostly worse than numeric output in open models. We hypothesize this may be due to the differences in post-training between the closed and open models, but we currently have no solid explanation for this difference, and we leave this for future work.
6 Conclusion and Future Work
We set out to discover if, or to what extent, autoregressive LLMs possess linguistic and other knowledge required to estimate the thematic fit of event participants. The tested LLMs set new state-of-the-art on each test set (and tied on Pado). Still, our takeaway is that thematic fit is not yet “solved”, in spite of this clear progress. Closed models [^11] We observed that GPT4.1 did worse than GPT4-Turbo, in spite of being newer, likely because 4.1 is too ‘cautious’ (§5). yielded higher results than open models, indicating that closed models, especially GPT4-Turbo, acquired much of the linguistic knowledge required for thematic fit estimation. However, no single axis setting consistently yielded the highest scores. Moreover, open models where better at filtering “bad” generated sentences, so we caution against rushing to interpret the results as showing that open models gained less relevant knowledge. It is still unclear if models need additional knowledge, or if the right elicitation has not been found yet.
We tested three axes for eliciting this knowledge: (1) Reasoning Form: adding Step-by-Step Prompting, (2) Input Form: adding generated sentences as context, and (3) Output Form: predefined categories vs. numeric scores.
(1) Surprisingly, Step-by-Step Prompting only helped closed models; Our analysis showed that invalid or irrelevant reasoning not only was unhelpful but actually hurt performance. This happened more in open models.
(2) Generated sentences hurt closed models; We saw LLMs struggle with generating sentences where the given argument is a better fit to a semantic role different than the given one. To tackle this, we used novel semantic filtering, which, to our surprise, worked better for open models, because the latter were better at filtering out False Positives.
(3) Predefined categorical output had a mixed effect (relative to numeric output): zero-to-small effect with closed models and more negative effect with open models.
In future work, we hope to improve Step-by-Step Prompting, following our analysis, especially by generating better sentences as context, and better filtering out “bad” sentences. If successful, LLMs may also be used to study issues in thematic fit, in line with futrell2025how, BabyLM efforts [^12] https://babylm.github.io, and others who believe that LLMs can be useful tools to study linguistic phenomena.
Limitations
Our experiments were evaluated on small datasets (a known limitation of this computational psycholinguistics subfield, as opposed to typical machine learning), so findings and conclusions should be assessed with caution. Our study considered only few LLMs, so one should be cautious about drawing conclusions for all LLMs.
LLMs can be very sensitive to prompt phrasing, so it is possible that different prompts may yield very different results. Also, prompt engineering is tricky; thus, tweaking the prompts may lead to different results. The richness of prompt variation (prompt engineering) depends on human resources (different people use language differently) and funding/time resources, all of which were limited. Future work may achieve better results with more resources, if available. Due to the scarcity of evaluation data, we could not use prompt optimization methods such as DSPy Khattab2023DSPyCD or fine-tune the models.
Last, all findings and conclusions here rely on data and annotations in English, as it is by far the most resource-rich. But thematic fit (and semantics in general) are thought of as universal, and of course relevant to all languages. Still, we would advise to be cautious of drawing universal conclusions before validating our findings on an additional, diverse set of languages.
Appendices
Appendix A Prompt Design
A.1 Simple Prompting with Head Lemma Tuples (Exp.1.1 — Exp.1.2)
[…]
  • Given the predicate ‘eat’, how well does the argument ‘pizza’ fit the PropBank Arg1 role?
    Reply only with a valid JSON dictionary {"Score": numeric_value} where numeric_value is a float number from 0 to 1. Avoid adding any text outside this JSON dictionary.
    Reply only with a valid JSON dictionary {"Score": string_value} where string_value that is one of "Perfect", "High", "Medium", "Low" or "Unlikely/Impossible". Avoid adding any text outside this JSON dictionary.
A.2 Semantic Filtering Process for Generated Sentences
Before executing experiments 2 and 4, the model is asked to generate $k$ diverse semantically coherent sentences containing an argument in a specific semantic role, given a predicate. Here we use $k=5$.
  • Give me five sentences with the predicate ‘eat’ and argument ‘pizza’ in the PropBank Arg1 role. Make the sentences as diverse as possible, while using the predicate as the main verb and keeping the given argument in the specified role. Make sure that each sentence is semantically coherent and the argument is in the given role.
LLMs may often generate a semantically incoherent or incompatible sentence with respect to the given role for the argument. Thus, we design a semantic filtering process of 4 waterfall Yes/No Steps. If the sentence passed the whole waterfall at least once out of three trials, it is considered semantically acceptable. Here are the prompts we used:
  • (2a) Here’s a sentence: ‘I ate a pizza with my friends’. Is the given sentence semantically coherent?
  • (2b) For the sentence: ‘I ate a pizza with my friends’. Does the sentence contain the predicate ‘eat’?
  • (2c) For the sentence: ‘I ate a pizza with my friends’. Does the sentence contain ‘pizza’ as an argument of the predicate ‘eat’?
  • (2d) For the same sentence: ‘I ate a pizza with my friends’. Given the definitions, instructions, and answers above, is ‘pizza’ in the PropBank Arg1 role for the predicate ‘eat’?
A.3 Simple Prompting with Generated Sentences (Exp.2.1 — Exp.2.2)
  • Given the following sentence, ‘I ate a pizza with my friends’, how well does the argument ‘pizza’ of the predicate ‘eat’ fit the PropBank Arg1 role?
A.4 Step-by-Step Prompting with Lemma Tuples (Exp.3.1 — Exp.3.2)
Experiments 3.x are the enhanced version of the baseline Exp.1.x, where the reasoning steps precede the final semantic-fitness prompt. The reasoning phase is done through Step-by-Step Prompting, where a series of four prompts helps reveal the reasons/internal knowledge that the model relies on to rate the thematic fit of the given data. These four prompts are:
  • What salient essential properties or core characteristics best describe the essence of ‘pizza’? [^13] In an early version of this experiment, only the semantic role name was mentioned. After analyzing the early version, we suspected that the LLMs may have not made the connection between roles such as ‘Arg1’ and their PropBank sense in this context. Therefore. We added the word ‘PropBank’ before the semantic role in the Pado and McRae datasets, as their semantic roles were PropBank-based, unlike the Ferretti datasets. This indeed improved the results on these sets by 2-3% (absolute).
  • Given the predicate ‘eat’, and the argument of it in the PropBank Arg1 role, what salient essential properties or core characteristics are most expected for this argument, to serve well in the PropBank Arg1 role?
  • Given the predicate ‘eat’, what PropBank roles are most suitable or likely for its argument ‘pizza’?
  • Given the predicate ‘eat’ with the argument ‘pizza’ in the PropBank Arg1 role, what essential properties does the argument have, properties which fit well its PropBank Arg1 role? And what essential properties this argument should have had (but are missing) in order for it to serve well in its PropBank Arg1 role?
A.5 Step-by-Step Prompting with Generated Sentences (Exp.4.1—Exp.4.2)
This last set of experiments combines Step-by-Step Prompting in Exp.3.x and the generated sentence input type as in Exp.2.x. In the later version of this experiment, the first four prompts were as in Exp.3.x but with the addition of the generated sentence in each prompt. Then the simple thematic fit prompt follows, also with one of the generated sentences as context, as in Exp.2.x :
  • Here’s a sentence: ‘I ate a pizza with my friends.’, What salient essential properties or core characteristics best describe the essence of ‘pizza’?
Appendix B Experimental Results with Variance
Table 7 and 8 show the Spearman correlation scores along with the variance for each experiment across the board. These tables support our explanation for the difference in performance between GPT4-Turbo and GPT4.1.
Model Dataset Fer-Loc Fer-Ins Pado McRae
$\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$
BG: GSD2015 .29 - .42 - .53 - .36 -
B0: NN RF .44 - .45 - .52 - .38 -
B1: ResRoFA-MT .46 - .48 - .53 - .43 -
B2: 20% subset v2 - - - - .43 - .44 -
B3a: Glove-shared-not-tuned .30 - .34 - .49 - .31 -
B3b: Glove-shared-tuned .29 - .29 - .53 - .33 -
B3c: Glove-shared-tuned with 20%M .35 - .43 - .59 - .43 -
Exp.1.1 .63 .121 .64 .132 .35 .168 .54 .159
Exp.1.2 .61 .081 .65 .120 .40 .142 .58 .131
Exp.2.1 .48 .085 .49 .075 .42 .149 .55 .136
Exp.2.2 .46 .063 .54 .081 .45 .154 .56 .130
Exp.3.1 .69 .097 .70 .115 .48 .125 .65 .137
Exp.3.2 .68 .108 .68 .132 .46 .118 .66 .138
Exp.4.1 .57 .077 .66 .074 .58 .130 .58 .133
Exp.4.2 .58 .091 .66 .092 .59 .132 .59 .145
Table 7: GPT4-Turbo Spearman Rank Correlation Coefficient ($\rho$) with variance.
Model Dataset Fer-Loc Fer-Ins Pado McRae
$\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$ $\rho$ $\mathbf{Var.}$
BG: GSD2015 .29 - .42 - .53 - .36 -
B0: NN RF .44 - .45 - .52 - .38 -
B1: ResRoFA-MT .46 - .48 - .53 - .43 -
B2: 20% subset v2 - - - - .43 - .44 -
B3a: Glove-shared-not-tuned .30 - .34 - .49 - .31 -
B3b: Glove-shared-tuned .29 - .29 - .53 - .33 -
B3c: Glove-shared-tuned with 20%M .35 - .43 - .59 - .43 -
Exp.1.1 .51 .046 .56 .105 .40 .145 .64 .158
Exp.1.2 .50 .051 .56 .105 .40 .147 .58 .149
Exp.2.1 .32 .024 .33 .050 .33 .096 .40 .058
Exp.2.2 .26 .027 .37 .061 .33 .114 .37 .066
Exp.3.1 .59 .042 .59 .082 .45 .111 .64 .118
Exp.3.2 .56 .062 .59 .097 .42 .127 .63 .136
Exp.4.1 .56 .030 .53 .056 .42 .104 .44 .061
Exp.4.2 .53 .043 .53 .071 .42 .121 .46 .074
Table 8: GPT4.1 Spearman Rank Correlation Coefficient ($\rho$) with variance.
Appendix C Earlier round of Experiment
A previous round of these experiments has been done on GPT4-Turboand CodeLlama2 (codeLlama2-70B-instruct) (roziere2023code) with a previous version of the prompts we use. The results of the earlier round are shown in Table 9 and Table 10.